MOM Summary 11/01/10 Due to the Inclement Weather it has perhaps been quieter than usual Water - Tank has been holding water fine. Some ice has been forming.
Published byModified over 5 years ago
Presentation on theme: "MOM Summary 11/01/10 Due to the Inclement Weather it has perhaps been quieter than usual Water - Tank has been holding water fine. Some ice has been forming."— Presentation transcript:
MOM Summary 11/01/10 Due to the Inclement Weather it has perhaps been quieter than usual Water - Tank has been holding water fine. Some ice has been forming in the tank. - Stopped going on the roof after the 7 th as conditions are too treacherous. -Display of water depth just outside MLCR seems fine. Could do with getting the conductivity readings fed back similarly. I believe A.Nichols has approached J. Governs wrt this. -Magnet Chiller controller outside of MLCR is crashing on a daily basis. Needs resetting every morning. Probable cause – Has been misbehaving since a component was removed. – Willie is also chasing this with John Governs. This box is currently only used for monitoring purposes.
MOM Summary 11/01/10 MICE HALL Decay Solenoid A larger Transfer Line collar is to be fitted on DS later today by WS and MH. This is to enable several adapters to be removed and will hopefully help improve the vacuum. Cryo Compressers A batch of fittings for the cryo compressors has arrived on the RAL site on Xmas Eve but have been mis-delivered. These are currently being chased by A.Nichols Other Work Willies team are currently drilling holes in the steelwork. Preparation work for the tracker solenoids. Linde Plant Servicing to run from 18 th to 22 nd Hall Survey completed this weekend. TOF1 slightly out. Willie sent results out.
MOM Summary 11/01/10 Daresbury - Guys are due to come Tue/Wed this week. -Installing Luminosity Monitors in target Synch - Replace component in target PSU - Install new interlock for target frame (will not be functional yet though) - Ensure that the target tip TC HW is correctly installed.
MOM Summary 11/01/10 Luminosity Monitors -Willie seems to have all of this in hand! -RAMS have all been sent off to Jane Vickers -P. Soler’s boys are due down today to bring the actual monitors. Paul is coming down himself on Tue or Wed -The stands for the monitors are in the MICE Hall and painted. Awaiting plate and lead blocks from G.Barber Wed – Pulling Cable from MLCR to synch. – Andy Nobbs / I.M. Job should be complete by the end of the week. DAQ - Spoke to JSG. He plans to come 26 th to 31 st of Jan to fix CAMAC. I have requested documentation so that a non-daq expert can plug leads back in.
MOM Summary 11/01/10 Running -No running planned for January because of shutdown -Running anticipated early in February Machine Physics (MP) 2 nd Feb? TBC -User Run Starts 16 th Feb – Confirmed -There will be a ISIS planning meeting later this month so that MICE can request beam time during MP. I will send out a request for what we would like to study. -UPDATE 14:00 -Alain has been discussing what we might like to do. I will draw up a list and circulate shortly